OpenStudy (anonymous):

how do i find the fourth roots of 256? I have an idea of how to do it but im confused with how to start

OpenStudy (mathteacher1729):

You're looking for a number "x" such that when raise it to the fourth power, you get 256. x^4 = 256 x*x*x*x = 256 Fortunately, you can find this by just starting to take small numbers to the fourth power. :) 1^4 = 1 2^4 = 2*2*2*2 = 16 3^4 = 3*3*3*3 = 81 4^4 = ... etc Hope this helps. :)

OpenStudy (anonymous):

I think you should find the sqroot 256 and then sqroot of the sqroot256

OpenStudy (mathteacher1729):

@PutuDani presents a very quick solution.

OpenStudy (anonymous):

u could prime factorize 256 and then that no. out which appears 4 times in the factorizations . since 256 = 2^8 u are going to take 2 out 2 times. so 2*2 = 4.